### v2

# Sub Specie Aeternitatis: Fourier Transforms from the Theory of Heat to Musical Signals

Abstract.
J. B. Fourier in his Théorie Analytique de la Chaleur of 1822 introduced, amongst other things, two ideas that have made a fundamental impact in fields as diverse as Mathematical Physics, Electrical Engineering, Computer Science, and Music. The first one of these, a method to find the coefficients for a trigonometric series describing an arbitrary function, was very early on picked up by G. Ohm and H. Helmholtz as the foundation for a theory of musical tones. The second one, which is described by Fourier’s double integral, became the basis for treating certain kinds of infinity in discontinuous functions, as shown by A. De Morgan in his 1842 The Differential and Integral Calculus. Both make up the fundamental basis for what is now commonly known as the Fourier theorem. With the help of P. A. M. Dirac’s insights into the nature of these infinities, we can have a compact description of the frequency spectrum of a function of time, or conversely of a waveform corresponding to a given function of frequency. This paper, using solely primary sources, takes us from the physics of heat propagation to the modern theory of musical signals. It concludes with some considerations on the inherent duality of time and frequency emerging from Fourier’s theorem.
Dramatis Personae
$X(f)=\int_{-\infty}^{\infty}x(t)e^{-i2\pi ft}dt$…………………………………………………….the Fourier transform,
$x(t)=\int_{-\infty}^{\infty}X(f)e^{i2\pi ft}df$………………………………………………………………..and its inverse,
$x(t)=\sum_{k=-\infty}^{\infty}X(kF)e^{i2\pi kFt}$………………………………………………………the Fourier series,
$X(f)=\sum_{n=-\infty}^{\infty}x(nT)e^{-i2\pi fnT}$………………………….the Discrete-time Fourier transform,
$X(k)=\sum_{n=0}^{N-1}x(n)e^{-i2\pi k\frac{n}{N}}$……………………………………….the Discrete Fourier transform,
$x(n)=\sum_{k=0}^{N-1}X(k)e^{i2\pi k\frac{n}{N}}$……………………………………………………………….and its inverse.
Act I
in which we see the characters’ in their formative years
1. Fourier
In Chapter III of his Théorie Analytique de la Chaleur [4, 5] from 1822, J. B. Fourier introduces the definition of the series that bears his name. He develops its formulation as two series of cosines and sines, which are shown to be independent of each other, one representing even functions and the other odd functions, respectively. The discussion culminates in §234, where we see the enunciation of the equation1
(1) $rf(x)=$ $\frac{1}{2}\int_{-r}^{+r}f(x)dx$
$+\cos\frac{\pi x}{r}\int f(x)\cos\frac{\pi x}{r}dx+\cos\frac{2\pi x}{r}\int f(x)\cos\frac{2\pi x}{r}dx+etc.$
$+\sin\frac{\pi x}{r}\int f(x)\sin\frac{\pi x}{r}dx+\sin\frac{2\pi x}{r}\int f(x)\sin\frac{2\pi x}{r}dx+etc.$
marked as Eq.(P) in his manuscript. This is followed by the two complementary half-series,
(2) $\frac{1}{2}rf(x)=$ $\frac{1}{2}\int_{0}^{+r}f(x)dx$
$+\cos\frac{\pi x}{r}\int f(x)\cos\frac{\pi x}{r}dx+\cos\frac{2\pi x}{r}\int f(x)\cos\frac{2\pi x}{r}dx+etc.$
and
(3) $\frac{1}{2}rf(x)=$ $\sin\frac{\pi x}{r}\int_{0}^{+r}f(x)\sin\frac{\pi x}{r}dx$
$+\sin\frac{2\pi x}{r}\int f(x)\sin\frac{2\pi x}{r}dx+etc.,$
marked as (M) and (N), respectively. Fourier also shows that the limits of the integration can be taken from $x=0$ to $x=2r$, where the whole interval represented by an arbitrary quantity $X$, by replacing $r$ in the denominator of the arguments to each cosine and sine term in Eq.(1).
Next in §235, Fourier makes the statement that is at the foundation of his theory,
Il résulte de tout ce qui a été démonstré dans cette section, concernant le développement des fonctions en séries trigonométriques, que si l’on propose une fonction $f(x)$, dont la valeur est représentée das un intervalle déterminé, depuis $x=0$ jusqu’à $x=X$, par l’ordonnée d’une ligne courbe tracée arbitrairement on pourra toujours développer cette fonction en une série qui ne contiendra que les sinus, ou les cosinus de multiples impairs. On emploiera, pour connaître les termes des ces séries, les équations (M), (N), (P)2. [4, p.258]
Effectively, this somewhat bold statement tells us that we can have any arbitrary $f(x)$ in an arbitrary interval, and replace it by a trigonometric series such as the one in Eq.(1). In addition to this, Fourier also remarks that
  • The series shown are always convergent for any function $f$ of a real variable $x$.
  • The sum of two arbitrary functions $f(x)$ and $\phi(x)$ corresponds to a series where each coefficient is a product of the corresponding coefficients in the series of these two functions.
  • In Eq.(1), if $r$ becomes infinitely large, each term becomes an infinitely small integrand, and the series is transformed into a definite integral.
Another important point is also made, that Eq.(1) can be represented by
(4) $F(x)=\frac{1}{\pi}\int_{-\pi}^{\pi}F(\alpha)d\alpha\left\{\frac{1}{2}+\sum_{i=0}^{\infty}\cos i(x-\alpha)\right\},$
where
(5) $\frac{1}{2}+\sum\cos i(x-\alpha)$
is a function of $x$ and $\alpha$, which if multiplied by an arbitrary function $F(\alpha)$ and integrated from $\pi$ to $\pi$, it becomes $\pi F(x)$. That particular function that maps $\alpha$ to $x$ is central to Fourier’s theorem, as we will see later. For now, we should note that this was taken by P.G. Lejeune-Dirichlet some years later in 1829 and given the form [10]
(6) $\frac{1}{2}+\sum\cos i(x-\alpha)=\frac{\sin\left(i+\frac{1}{2}\right)(\alpha-x)}{2\sin\frac{1}{2}(\alpha-x)},$
which later became known as the Dirichlet kernel. We shall also revisit this later on in this paper.
2. Ohm and Helmholtz
The ideas developed by Fourier, embodied in Eq.(1), enter Music Theory through the work of H. Helmholtz as exposed in his On the Sensations of Tone [9], and founded on principles established by G. Ohm. In 1843, Ohm defends the notion that complex tones can be described by a trigonometric series, which, although already established earlier by D. Bernoulli, M. Poisson, and others, had been perceived by him to be under attack3. In order to do so, he brings out the Fourier series, now in the form we are accustomed to see [12, p.519],
(7) $F(t)=A_{0}+$ $A_{1}\cos\pi\frac{t}{l}+A_{2}\cos 2\pi\frac{t}{l}+A_{3}\cos 3\pi\frac{t}{l}+$.
$B_{1}\sin\pi\frac{t}{l}+B_{2}\sin 2\pi\frac{t}{l}+B_{3}\sin 3\pi\frac{t}{l}+$.,
with coefficients $A_{n}$ and $B_{n}$ now defined separately
(8) $A_{0}$ $=\frac{1}{2l}\int F(t)dt$
$A_{1}$ $=\frac{1}{l}\int F(t)\cos\pi\frac{t}{l}dt\quad B_{1}=\frac{1}{l}\int F(t)\sin\pi\frac{t}{l}dt$
$A_{2}$ $=\frac{1}{l}\int F(t)\cos 2\pi\frac{t}{l}dt\quad B_{2}=\frac{1}{l}\int F(t)\sin 2\pi\frac{t}{l}dt$
$A_{3}$ $=\frac{1}{l}\int F(t)\cos 3\pi\frac{t}{l}dt\quad B_{3}=\frac{1}{l}\int F(t)\sin 3\pi\frac{t}{d}tl$
$u.s.w\quad\quad\quad\quad\quad\quad\quad\quad u.s.w,$
and integrated from t=lt=-l to $t=l$.
Ohm states that the ear performs the analysis of complex tones into simple sinusoidal partials, and that the mathematical form is given by Eqs. (7) and (8). This is the basis on which Helmholtz establishes his theory. It is interesting to note that he appears first to cast some doubt on the validity of the mathematics involved,
The theorem of Fourier here adduced shews first that it is mathematically possible to consider a musical tone as a sum of simple tones, in the meaning we have attached to the words, and mathematicians have found it convenient to base their acoustic investigations on this mode of analysing vibrations. But it by no means follows that we are obliged to consider the matter in this way. We have rather to enquire, do these partial constituents of a musical tone such as the mathematical theory distinguishes and the ear perceives, really exist in the mass of air external to the ear? Is this means of analysing forms of vibration which Fourier’s theorem prescribes and renders possible, not merely a mathematical fiction, permissible for facilitating calculation, but not necessarily having any corresponding actual meaning in things themselves? What makes us hit upon pendular vibrations, and none other, as the simplest elements of all motions producing sound? [9, p.36]
However, he follows on to indicate that Fourier’s ideas have a probable meaning in nature because of the fact that the ear really effects the same analysis [9, p.36], which was what Ohm indicated in his paper4. Helmholtz also explores the idea that simple tones (or partials) have an existence outside of the ear through the demonstration of physical effects such as resonance and sympathetic vibrations. He produces what he claims as experimental proof of Ohm’s partial-tone law. Eventually, he places Fourier’s theorem as the basis of his theory of harmony,
Ultimately, then, the reason of the rational numerical relations of Pythagoras is to be found in the theorem of Fourier, and in one sense this theorem may be considered as the prime source of the theory of harmony5. [9, p.36]
We should also note that, implicit in Ohm’s and Helmholtz’s use of the Fourier series, is the assumption the resulting function of time $F(t)$ in Eq.(7) is periodic, repeating every $2l$ seconds. This is not actually something we find explicitly in Fourier’s theory. However, this was shown later in 1874 by J. O’Kinealy to be the case [13]. From Taylor’s theorem, a function of two variables may be expanded as
(9) $f(x+\lambda)=e^{\lambda\frac{\delta}{\delta x}}f(x).$
Therefore, for a periodic function $f(x+\lambda)=f(x)$, we have
(10) $\left(e^{\lambda\frac{\delta}{\delta x}}-1\right)f(x)=0,$
whose auxiliary equation roots are $e^{\lambda p}=1$, with $p=\delta/\delta x$.
[…]
(11) $f(x)=A_{0}$ $+a_{1}e^{i2\pi/\lambda}x+a_{2}e^{i4\pi/\lambda}x+$.
$+b_{1}e^{-i2\pi/\lambda}x+b_{2}e^{-i4\pi/\lambda}x+$.,
which is Eq.(7) expressed using complex exponentials. In this form we have
(12) $a_{n}$ $=\int f(x)e^{-i2n\pi x/\lambda}dx$
$b_{n}$ $=\int f(x)e^{i2n\pi x/\lambda}dx,$
integrated over 0 to $\lambda$, with $i$ now denoting $\sqrt{(}-1)$.
3. Fourier
The integral form of Fourier’s series alluded to earlier is probably the most general statement of his mathematical ideas, and is what is normally referred to as the Fourier theorem, sometimes also qualified as the double integral bearing his name. It is developed in Chapter IX, in a similar manner to how Eq.(1) was produced.
Taking two functions, $F(x)$ and $f(x)$, satisfying F(x)=F(x)F(x)=F(-x) and f(x)=f(x)f(x)=-f(-x), respectively. The two integrals,
(13) $\frac{1}{2}\pi F(x)=\int_{0}^{\infty}dq\cos qx\int_{0}^{\infty}dxF(x)\cos qx$
and
(14) $\frac{1}{2}\pi F(x)=\int_{0}^{\infty}dq\sin qx\int_{0}^{\infty}dxf(x)\sin qx$
are the half-transforms that can be used to arrive at $\phi(x)=F(x)+f(x)$. These are exactly the result of extending the range of summation in Eqs. (2) and (3) to infinity as indicated by Fourier earlier in Chapter III.
This is the basis of Fourier’s theorem, first given as
(15) $\pi\phi(x)=\int_{0}^{\infty}$ $dq\sin qx\int_{-\infty}^{\infty}d\alpha\phi(\alpha)\sin q\alpha$
$+$ $\int_{0}^{\infty}dq\cos qx\int_{-\infty}^{\infty}d\alpha\phi(\alpha)\cos q\alpha,$
which as can be seen introduces a separate variable $\alpha$ in the right-hand-side $\phi()$. We can view this as the input to the expression. The reason for this is that his aim is to develop an equation like Eq.(5) as an integral sum. This he does in the following, labelled (E) in § 361,
(16) $\phi(x)=\frac{1}{\pi}\int_{-\infty}^{+\infty}d\alpha\phi(\alpha)\int_{0}^{\infty}dq\cos q(x-\alpha),$
to which he adds the remark,
Ainsi la fonction représentée par l’integrale définie $\int dq\cos q(x-\alpha)$ a cette singulière propriété, que si on la multiplie par une fonction quelconque, $\phi(\alpha)$ et par $d\alpha$, et si l’on intègre par rapport à $\alpha$ entre des limites infinies, le résultat est égal à $\pi\phi(x)$; en sorte que l’effet de l’intégration est de changer $\alpha$ en $x$ et de multiplier par le nombre $\pi$.6 [4, p.449]
The significance of this statement in various applications became apparent as Fourier’s work was followed by others. An example of this was given by A. De Morgan, where Eq.(16) is used to describe any function varying continuously or discontinuously, in any manner whatever, from $x=$-$\infty$ to $x=\infty$. To show this, he starts by putting Fourier’s theorem in the form [11, p.628]
(17) $\phi(x)=\frac{1}{\pi}\int_{0}^{\infty}d\omega\int_{-\omega}^{+\omega}\cos\omega(\nu-x)\phi(\nu)d\nu.$
Following this, he sets up $\phi(\nu)$ as made of three segments, $\phi(\nu)=0$, from $\nu=$-$\infty$ to $\nu=a$; $\phi(\nu)=\psi(\nu)$, from $\nu=a$ to $\nu=b$; and $\phi(\nu)=0$, from $\nu=b$ to $\nu=\infty$. Now he replaces $\omega$ and $\omega$ by $a$ and $b$. Using the constant function $\psi(\nu)=1$, we have
(18) $\phi(x)=\frac{1}{\pi}\int_{0}^{\infty}d\omega\int_{a}^{b}\cos\omega(\nu-x)d\nu.$
With this, he arrives at
[…]
which describes what is now known as a rectangular function. We will later revisit this in another guise.
De Morgan follows this by showing that with Fourier’s theorem in the form
(20) $\frac{1}{\pi}\int_{0}^{\infty}d\omega\int_{a}^{b}\cos\omega(\nu-x)\phi(\nu)d\nu,$
which he labelled $F_{a}^{b}\phi(x)$, e.g. $F_{a}^{b}1$ for Eq.(19), we may describe any function in segments, $a$ to $b$, $b$ to $c$, $c$ to $d$, etc, as the sum
(21) $f(x)=F_{a}^{b}\phi(x)+F_{b}^{c}\psi(x)+F_{c}^{d}\chi(x)+etc.$
In this case, we have $f(a)=\frac{1}{2}\phi(a)$, $f(b)=\frac{1}{2}[\phi(b)+\psi(b)]$, $f(c)=\frac{1}{2}[\psi(c)+\chi(c)]$, etc.
De Morgan also provides another example where we have a ramp $f(x)=x$ from 0 to 1, zero elsewhere, and notes that the result is the limit of7
(22) $\frac{1}{\pi}\int_{0}^{\infty}e^{-k\omega}d\omega\int_{a}^{b}\cos\omega(\nu-x).\nu d\nu,$
as $k\rightarrow 0$, except for the discontinuity point at $x=1$, where it is $1/2$ [11, p.628].
We should remark that this application of Fourier’s theory is quite distinct from that of Ohm and Helmholtz. Here, instead of exploring periodic functions, the focus is on the finding analytical expressions for discontinuous cases. To bridge these two perspectives, we can say that this work gives us the means of understanding, for example, a waveform that is turned on and off, a single cycle that is not repeated, etc. This has implications for, amongst other things, our understanding of time and frequency in the context of the Fourier transform, and the possibility of slicing functions at will to manipulate them. We shall return to this later in the paper.
4. Dirac
Further tools for handling these emerging questions about infinity are provided by P.A.M. Dirac’s delta. A notation for this was first given in his 1927 paper8 The Physical Interpretation of the Quantum Dynamics [2], where he introduces the symbol $\delta(x)$,
(23) $\delta(x)=0\quad\textrm{when}\quad x\neq 0$
and9
(24) $\int_{-\infty}^{\infty}\delta(x)dx=1.$
Dirac states at the outset that $\delta$ is not a proper function of $x$, and we shall continue to observe this by omitting the word function in relation to it. He also remarks that it may be seen as a limit of a certain sequence of functions. The main idea of this representation is to give us a form to treat a mathematical object that is zero everywhere except at an infinitely small area around 0.
There are a number of important properties that are relevant in the context of this paper. The first one of these is
(25) $\int_{-\infty}^{\infty}f(x)\delta(x)dx=f(0),$
for any continuous $f(x)$. From the earlier definitions, it seems clear that the product by $\delta(x)$ and integration effectively select the value of the function for $x=0$.
We can now deduce
(26) $\int_{-\infty}^{\infty}f(x)\delta(x-a)dx=f(a)$
with $a$ standing for any real number. Dirac makes the following remark in relation to this equation:
the process of multiplying a function of $x$ by $\delta(x-a)$ and integrating over all $x$ is equivalent to substituting $a$ for $x$. [3, p.59]
This brings us back to Fourier’s theorem, where a similar statement was made about the definite integral in Eq.(16). Indeed, we can now spot that
(27) $\int_{0}^{\infty}dq\cos q(x-\alpha)=\pi\delta(x-\alpha),$
A proof is fairly straightforward:
  1. (1)
    For the case where $x\neq\alpha$, the left hand-side of Eq.(27) vanishes10, so we have $\delta(x)=0$ for $x\neq 0$.
  2. (2)
    With $x=\alpha$, we can revert to Eq.(16), where we have
    (28) $\frac{1}{\pi}\int_{-\infty}^{+\infty}\phi(x)\left\{\int_{0}^{\infty}\cos q(0)dq\right\}dx=\phi(0),$
    which can be compared to Eq.(25) to prove $\delta(x)$ for $x=0$ and $\pi^{-1}\int\cos q(x)dq=\delta(x)$.
With this, the connection between Fourier’s theory and the infinities described by the $\delta$ are made evident. Dirac sees the $\delta$ as an improper function without a well-defined value. However, when appearing as a factor in an integrand, it leads to the integral having a well-defined value. Therefore its appearance should not be regarded as a lack of rigour in this case, but as a convenient notation.
Two final aspects of the $\delta$ are worth noting as they have relevance to the matter at hand. Firstly, the $\delta$ can also be defined as the derivative of the step function, originally introduced by O. Heaviside [8, p.510], and quoted by Dirac as11
(29) $\epsilon(x)=\begin{cases}0,&\quad x<0\\ 1,&\quad x>0.\end{cases}$
and as proof $\epsilon^{\prime}(x)$ can be replaced for $\delta(x)$ in Eq.(25) and it can be solved to get as a result f(0)f(0)12. This demonstrates that a delta is involved whenever we try to differentiate a discontinuity.
The second point is that $\delta(x)$ can be differentiated $n$ times, resulting in $\delta^{\prime}(x)$, $\delta^{\prime\prime}(x)$, and so on. Associated with this we have the fact that while, trivially, $\delta(x)=\delta(-x)$, therefore even, its derivative is odd, $\delta^{\prime}(x)=-\delta^{\prime}(-x)$. Following this, we can also deduce that the second-order derivative is even, the third is odd, and so on. Any derivative of order $n$, $\delta^{(n)}(x)$ can replace it in Eq.(26) producing as a result the respective derivative of $f(a)$, $f^{(n)}(a)$.
Act II
in which the characters operate in time and frequency
5. Transforms
We have now effectively all the tools to explore the Fourier transforms, giving it a physical interpretation that was mostly missing so far. From Eq.(15), we can develop a modern musically-meaningful form of his eponymous Transform. We will proceed to do this by
  • splitting the single equation into two,
  • replacing the real-valued sinusoids by complex exponentials; and
  • employing functions of time $t$ in seconds and frequency $f$ in Hz.
This gives the pair
(30) $X(f)=\int_{-\infty}^{\infty}x(t)e^{-i2\pi ft}dt,$
and
(31) $x(t)=\int_{-\infty}^{\infty}X(f)e^{i2\pi ft}df.$
In this way, the transforms are set in a very compact form, and it is possible to observe the relationship between time on side and frequency on the other. We can use the term spectrum as introduced by D. Gabor [6] to denote the frequency domain, and waveform for the time domain. It is also possible to develop an intuition about the operation of the transforms. Firstly, we should interpret the complex exponential for what it is in this context, a sinusoid with arbitrary phase and frequency. Secondly, we can observe that, starting with a waveform defined by $x(t)$, we use a product of it with a sinusoid of frequency $f$, and by integrating over all $x$ we get a complex number that tells us the amplitude and phase of the waveform at that frequency. We may even go as far as using a radio receiver metaphor, the frequency $f$ is what we set in our dial, and if there is anything there in the waveform, we will get a result.
From an inverse perspective, once we have $X(f)$ defining amplitudes and phases in the spectrum, we can recompose the waveform by multiplying these with complex sinusoids at all frequencies $f$, then effecting an integral sum of all these partial waveforms. This can be described as a form of additive synthesis. Additionally, having the spectrum $X(f)$, allows us to pick individual partials if we wish, instead of applying the full summation.
6. Zero Frequency and Zero Time
With that in mind, we can now exact the mathematical form in which this can be represented. Starting with the simplest of the waveforms, one which is just a constant, $x(t)=1$ for all $x$, we have [3, p.95]
(32) $X(f)=\int_{-\infty}^{\infty}e^{-i2\pi ft}dt=\delta(f).$
This result can be verified in various ways, intuitively first by noting that both imaginary and real parts of this equation vanish for $x\neq 0$, and at $X(0)$ the limit of the sum is 1. Secondly, we can replace it in Eq.(25), and solve it to get exactly $X(0)$, and finally, we can use Fourier’s theorem and Eq.(28)13.
The physical interpretation is as follows, since for $f=0$, we have $\cos(2\pi ft)=1$ (and $\sin(2\pi ft)=0$), this is a signal with energy concentrated in one single frequency, 0Hz. We commonly call this, with reference to electrical signals, a direct current (DC) waveform. We should note the exact role of the delta in this context, as a representation of the fact that waveform amplitude is fully concentrated at a single point in the frequency axis.
Conversely, let’s consider a signal that is concentrated at one point in time, and nowhere else, an infinitesimally small pulse. This can be represented by $x(t)=\delta(t)$, and we should very well expect a similar result by employing the inverse transform,
(33) $x(t)=\int_{-\infty}^{\infty}e^{i2\pi ft}df=\delta(t).$
That is, if the energy of a waveform is concentrated at a single time point, then it is unity everywhere in the frequency domain, $X(f)=1$. These two results should be regarded as the foundation for handling the duality of time and frequency. Another way of putting this is to say that if a musical signal is defined precisely in time, it cannot be determined in frequency; conversely if it is exactly defined in frequency, it cannot be determined in time. This result is of course a fundamental fact in Quantum Physics in the form given by Heisenberg’s principle of uncertainty [3, p.98], as well as in the theory of musical signals as interpreted by Gabor [6]14.
7. Sinusoids
Now we can extend these results sideways, so to speak, taking advantage of Eq.(26). Let’s consider a real-valued sinusoid with amplitude $A$, frequency $F$, and phase $\theta$, which can be described by
(34) $x(t)=A\cos(2\pi Ft+\theta)=A[a\cos(2\pi Ft)-b\sin(2\pi Ft)]$
Applying the transform, and using Eqs. (32) and (26), we get
(35) $X(f)$ $=\int_{-\infty}^{\infty}A[a\cos(2\pi Ft)-b\sin(2\pi Ft)]e^{-i2\pi ft}dt$
$=\frac{A}{2}\left[a\{\delta(f-F)+\delta(f+F)\}+ib\{\delta(f-F)-\delta(f+F)\}\right],$
that is, energy concentrated at two points in the frequency axis, at $f$ and f-f, each with half amplitude15. Notice that the result is generally complex, and the values of $a$ and $b$ determine the phase of the sinusoid as per the usual formula for converting rectangular to polar forms of complex numbers. The spectra of real-valued waveforms is always split into negative and frequency sides, which are symmetric around 0 Hz for their real part, and anti-symmetric for their imaginary part. This type of symmetry is called Hermitian [3, p.68]. From this result, we can deduce that if the waveform is made up of a mixture of various sinusoids, its spectrum will be a sum of deltas at each absolute frequency, with half of their amplitudes on the positive side and half on the negative. We should expect it to be zero elsewhere16.
The converse case is also interesting, what is the time domain function corresponding to a spectrum consisting of a single complex exponential, $X(f)=e^{-2\pi\tau f}$? We can proceed along similar lines to obtain
(36) $x(t)=\int_{-\infty}^{\infty}e^{2\pi(t-\tau)f}df=\delta(t-\tau),$
which can be given the physical interpretation of a time delay. This makes the pulse appear at time $t=T$ rather than at $t=0$. In fact, a relevant notion that arises from this, intuitively, is that if we take the product of an arbitrary spectrum and a complex exponential such as this, we will be applying a time delay of sorts to a waveform.
8. Series
What was said about the energy concentration in a waveform that consists solely of sinusoids can now be extended to give a generic formula for the Fourier series. If we have a waveform composed of sinusoids at exact multiples of a frequency $F$, then a discrete sum can be used to reconstitute it from its spectrum $X(f)$,
(37) $x(t)=\sum_{k=-\infty}^{\infty}X(kF)e^{i2\pi kFt},$
since we know this is composed purely of a sum of $\delta(f\pm kF)$. Moreover, because of the earlier results in §2, we should expect that the waveform is periodic over a period $T=\frac{1}{F}$ and so the spectrum only needs to be computed over this interval, for discrete frequencies $nF$,
(38) $X_{F}(n)=\int_{-\frac{T}{2}}^{\frac{T}{2}}x(t)e^{-i2\pi nFt}dt,$
which certainly simplifies things. This pair of transform-series computes a waveform from a discrete sequence of evenly-spaced frequency values.
Completing the set, we have a transform that takes an evenly-spaced discrete sequence of waveform values, and computes its continuous-frequency spectrum,
(39) $X(f)=\sum_{n=-\infty}^{\infty}x(nT)e^{-i2\pi fnT}$
for which, we should expect the inverse transform to be an integral over an interval $F=\frac{1}{T}$,
(40) $x_{T}(n)=\int_{-\frac{F}{2}}^{\frac{F}{2}}X(f)e^{i2\pi nfT}dt,$
producing a sequence of time values at evenly-spaced intervals.
These results point to the general principle that the spectrum of a discrete waveform is periodic in frequency, with period $F$ corresponding to the reciprocal of the time interval $T$ between points in the series. Conversely, the waveform computed from a discrete spectrum is also periodic, with period $T$, corresponding to the reciprocal of the frequency interval $F$ between points in the series. We shall return to this later as it gives the foundation for a fully discrete transform pair.
9. Dirac Comb
At this stage, we can expand our results by introducing another useful notation as an aid for more complicated cases. Reverting to the Dirichlet kernel introduced earlier in Eq.(6), if we rewrite it using complex exponentials, we have
(41) $x(t)=\sum_{n=-\infty}^{\infty}e^{2\pi nt}=\sum_{n=-\infty}^{\infty}\delta(t-n)=\Sha(t),$
which will be referred to as the Dirac comb. Similarly to the delta, we should refrain from it calling a function, and our use in this context will be simply as an operator in expressions (as demonstrated later). Heuristically, it can be interpreted as an impulse train.
An interesting property of the comb is this,
(42) $X(f)$ $=\int_{-\infty}^{\infty}\sum_{n=-\infty}^{\infty}\delta(t-n)e^{-2\pi tf}dt=\sum_{n=-\infty}^{\infty}e^{-i2\pi fn}$
$=\sum_{n=-\infty}^{\infty}\delta\left(f-n\right)=\operatorname{III}(f),$
which tells us that the spectrum of a comb waveform is also a comb, and conversely.
10. Convolution
It is useful now to determine what is the frequency domain function corresponding to the product of two waveforms. Starting with the expression $x(t)=g(t)h(t)$, and applying the forward transform of Eq. (30), we obtain
(43) $X(f)$ $=\int_{-\infty}^{\infty}x(t)e^{-i2\pi ft}dt=\int_{-\infty}^{\infty}g(t)h(t)e^{-i2\pi ft}dt$
$=\int_{-\infty}^{\infty}g(t)\int_{-\infty}^{\infty}H(\phi)e^{i2\pi\phi t}d\phi e^{-i2\pi ft}dt$
$=\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}H(\phi)g(t)e^{-i2\pi(f-\phi)t}dtd\phi=\int_{-\infty}^{\infty}H(\phi)G(f-\phi)d\phi,$
which introduces an operation called the convolution of $H(f)$ and $G(f)$. Well-known trigonometric identities such $\cos(\omega)\cos(\theta)=\frac{1}{2}[\cos(\omega-\theta)+\cos(\omega+\theta)]$ follow from this result. It should also not surprise us that the product of two spectra corresponds to the convolution of two waveforms, which can be demonstrated by the same methods applied here.
What is the physical interpretation of these results? We can observe that there are three separate operations wrapped in the convolution,
  1. (1)
    translation;
  2. (2)
    product; and
  3. (3)
    (integral) sum.
This should give us an intuition as to what is taking place here. We slide one function continuously for all $\phi$, taking the products of this and the other function at $f$, then we sum it all together. We can think of this as an accumulation of copies of one function, scaled by another at every infinitesimal point of the domain. Or simply put, a smear over frequency or over time. This becomes much simpler to picture if we apply it to discrete sequences, as the integral becomes just a sum.
11. Discrete Time and Discrete Frequency
We now have the means to go from continuous to discrete time in a more rigorous way. By reasoning that the process of forming a sampled17 discrete sequence of evenly-spaced values from a continuous function $x(t)$ can be equated to applying a set of deltas at each point, we can then apply a Dirac comb as an operator to represent this process18,
[…]
which may be simply referred to as sampling. The underscore $T$ now refers to the sampling interval.
Intuitively, from §10 we should expect the spectrum of this sequence to be a convolution of the spectra of $x(t)$ and the Dirac comb. This can be verified by
(45) $X_{T}(f)$ $=\int_{-\infty}^{\infty}x_{T}(t)e^{-2\pi ft}dt$
$=\int_{-\infty}^{\infty}X\left(f-\phi\right){}_{T^{-1}}(\phi)d\phi,$
where $T^{-1}$ is now known as the sampling frequency. It also confirms the general principle from §8, that the spectrum of a discrete time series is periodic in frequency.
The physical interpretation of this is as follows. A periodically sampled function of time has a spectrum that repeats at intervals of the sampling frequency $F$, which can be thought of as images of a base frequency band ranging from $\frac{F}{2}$ to $\frac{F}{2}$. This implies that we only need to look at this base band to know what the spectrum of a sampled waveform is.
Now, using the same principles, we can determine what happens when we sample a spectrum instead,
[…]
and reconstitute a waveform from it,
(47) $x_{F}(t)$ $=\int_{-\infty}^{\infty}X_{F}(f)e^{2\pi ft}dt$
$=\int_{-\infty}^{\infty}x\left(t-\tau\right)\operatorname{III}_{F^{-1}}(\tau)d\tau.$
This is effectively the result of the series discussed in §8 put in a more generic form. It tells us precisely that the sampled spectrum produces a periodic waveform with period $F^{-1}$. We should note that these equations do not imply that the underlying continuous-time waveform $x(t)$ has a spectrum that is band-limited to an interval $F$ or that the spectrum $X(f)$ corresponds to a waveform that is time-limited to an interval $T$. If they are not, there will be an overlap, in frequency or in time, between the repeated copies of the spectrum or waveform. This effect is called aliasing19
12. Finite Time and Finite Frequency
We should now turn our attention to defining what goes on when we turn a waveform on and off. This reverts back to the ideas in §3, where a focus was placed on this exact issue. We may equate the turning on/off of a waveform to the application of a function that is 1 for a given time interval and zero elsewhere, for a which an expression was derived in Eq.(19). With such interval defined evenly around the origin as $T$, this is commonly known as the rectangular function $\Pi_{T}(t)$. For $T=1$ we then have,
[…]
$=\frac{2}{\pi}\int_{0}^{\infty}\frac{\sin(\omega)\cos(\omega t)}{\omega}d\omega=\frac{1}{\pi}\int_{-\infty}^{\infty}\frac{\sin(\omega)}{\omega}e^{\omega t}d\omega=$
$=\int_{-\infty}^{\infty}\frac{\sin(\pi f)}{\pi f}e^{2\pi ft}df=\int_{-\infty}^{\infty}\operatorname{sinc}(f)e^{2\pi ft}df,$
which tells us that its spectrum is a sinc function. Similarly, it would not be again a surprise to see that the time domain waveform corresponding to a rectangular spectrum is also a sinc function.
We can now apply this to our problem of switching on/off a waveform20,
(49) $x_{T}(t)=x(t)\Pi_{T}(t)$
which corresponds to the spectrum21
(50) $X_{T}(f)$ $=\int_{-\infty}^{\infty}x_{T}(t)e^{-2\pi ft}dt$
$=\int_{-\infty}^{\infty}X\left(f-\phi\right)\operatorname{sinc}_{T}(\phi)d\phi$
Conversely, we can also cut the spectrum in a given band using the rectangular function,
(51) $X_{F}(t)=X(f)\Pi_{F}(f)$
resulting in the waveform
(52) $x_{F}(t)$ $=\int_{-\infty}^{\infty}X_{F}(f)e^{2\pi ft}dt$
$=\int_{-\infty}^{\infty}x\left(t-\tau\right)\operatorname{sinc}_{F}(\tau)d\tau$
The physical interpretation of this allows us, amongst other things, to treat the effects of aliasing described earlier. If we take a continuous waveform and apply the convolution to a sinc function, the resulting spectrum is the product of a rectangular function and the waveform spectrum. This therefore eliminates all frequencies outside a given base band. When we sample the resulting waveform we will be sure that no frequency aliasing effects are due. Conversely, if we take the product of a rectangle with width $T$ and a given waveform, the resulting spectrum is the convolution of a sinc function and the spectrum of the waveform. We can then sample that spectrum at multiples of $F=\frac{1}{T}$ without any concern for time aliasing.
To give a complete picture of our current state of affairs, we can combine the operations of sampling and cutting waveforms in one single expression,
(53) $x_{TS}(t)=x(t){}_{T}(t)\Pi_{S}(t).$
This corresponds to the continuous spectrum
(54) $X_{TS}(f)$ $=\int_{-\infty}^{\infty}x_{TS}(t)e^{-2\pi ft}dt$
$=\int_{-\infty}^{\infty}X_{T}\left(f-\phi\right)\operatorname{sinc}_{S}(\phi)d\phi,$
The sampled, band-limited spectrum defined as
(55) $X_{FS}(f)=X(f){}_{F}(f)\Pi_{S}(f),$
produces the waveform
(56) $x_{FS}(t)$ $=\int_{-\infty}^{\infty}X_{FS}(f)e^{2\pi ft}df$
$=\int_{-\infty}^{\infty}x_{F}\left(t-\tau\right)\operatorname{sinc}_{S}(\tau)d\tau$
13. Discrete Transforms
If we can cut a slice of a sampled waveform as a sequence of $N$ values, then using the ideas developed earlier, we can obtain a spectrum which consists of $N$ frequency values, each one a complex number. Here is how we put it,
(57) $X(k)=\sum_{n=0}^{N-1}x(n)e^{-i2\pi k\frac{n}{N}},$
for $k\in\mathbb{Z}$. This completely describes the discrete spectrum, which is periodic, with period $N$. The inverse transform, starting with a discrete spectrum $X(k)$ with $N$ points, can be put as
(58) $x(n)=\sum_{k=0}^{N-1}X(k)e^{i2\pi k\frac{n}{N}}$
for $n\in\mathbb{Z}$. The two transforms can be used transparently, e.g. forwards and inverse, like the continuous versions, to recover the original sequence, but a scaling factor $1/N$ needs to be applied once in either direction22.
In order for us to link $k$ to frequencies $f$ in Hz, we first set $t=nT$, where $T$ is the sampling period, so the time series is set at 0,T,2T,,(N1)T0,T,2T,...,(N-1)T seconds. Thus $F=\frac{1}{T}$ is our sampling frequency. Then the frequency points $k$ are set at 0,F/N,2F/N,,(N1)F/N0,F/N,2F/N,...,(N-1)F/N Hz. However, due to the periodicity in frequency, the points from $N/2$ to N1N-1 also refer to the frequencies F/2-F/2 to F/N-F/N, the negative side of the base band spectrum. It is more convenient to read the second half of the discrete transform spectrum as referring to negative frequencies, rather than their positive images.
The discrete transform may have two interpretations with regards to the underlying continuous-time functions, both equally valid. From the perspective of the forward transform, the first one is that we are looking at a waveform that is zero for all $t$, except for a time window of $NT$ seconds. In this case the underlying continuous-frequency spectrum is the convolution of a sinc function and the sampled spectrum. This latter function can be represented in continuous time by a sum of scaled deltas at each sampling point $\frac{kF}{N}$. This is what Eq.(54) gives as a result, and is equivalent to the methods explored in §3.
The second interpretation is that the waveform is a single cycle of a periodic function of time, which repeats from $t=$-$\infty$ to $t=\infty$, eternally. In this case the underlying continuous-time spectrum is zero everywhere, except at the sampling points, where it is effectively the same as the sampled spectrum. It is a sum of scaled deltas spaced evenly in frequency. While both are equally good representations, the first one, in conjunction with the principles from §3, allows us to employ the discrete transforms to successive blocks of a waveform successfully to obtain its spectrum. The second one allows us to create single-cycle waveforms in oscillators for wavetable sound synthesis.
Similar points can be made to the underlying continuous waveform of a sampled spectrum, if we take the spectrum as periodic, with period $F$, stretching from $f=$-$\infty$ to $f=\infty$, then the underlying time-domain waveform is exactly the same as a a sequence of scaled deltas at the sampling points. That is, infinitely small areas with zeros in between. On the other hand, if we take it as fully band-limited to $\frac{F}{2}$ to $\frac{F}{2}$, with $X(f)=0$ elsewhere, then the reconstruction of the continuous-time signal from samples is a convolution of sinc functions by these scaled deltas, that is, a smooth waveform.
In both time and frequency-domain cases, whenever we are taking a slice of the function within a given interval, we are multiplying it by a rectangular function. The result, as we have seen, is a convolution of the corresponding waveform or spectrum with a sinc function. Without this, we are limited to a single interpretation. This tells us that the continuous-time functions in both domains are a sum of translated, scaled deltas, periodic, and stretching infinitely.
Epilogue
the show must go on…
Fourier’s theory and its refinements that followed are unquestionably a success. As far as the theory of musical signals goes, they represent a great achievement in elucidating the nature of sounds and, more recently, their transformations. However, in order for it to be more generally applicable, it must be further developed. Certain types of musical signals are not so well described by the current state of affairs; for example, what is the representation of a musical signal where changes occur, such as for instance, a mix of notes appearing and disappearing, glissandos, noise-like sounds, etc? That surely has a spectrum, but what is it? J. Ville puts the question very elegantly,
Si nous considérons en effet un morceau contenant plusieurs mesures (ce qui est le moins qu’on puisse demander) et qu’une note, la par exemple, figure une fois dans le morceau, l’analyse harmonique nous présentera la fréquence correspondante avec une certaine amplitude et une certaine phase, sans localiser le la dans le temps. Or, il est évident qu’au cours du morceau il est des instants où l’on n’entend pas le la. La représentation est néanmoins mathématiquement correcte, parce que les phases des notes voisines du la sont agencées de manière à détruire cette note par interférence lorsqu’on ne l’entend pas et à la renforcer, également par interférence, lorsqu’on l’entend; mais s’il y a dans cette conception une habileté qui honore l’analyse mathématique, il ne faut pas se dissimuler qu’il y a également une défiguration de la réalité: en effet, quand on n’entend pas le la, la raison véritable est que le la n’est pas émis23. [14, p.64]
The difficulties we face do not necessarily arise as a failure of the theory, but from a fact of reality, to which we have alluded earlier. We may say that on one hand it is not possible to define frequency without reference to time, and on the other, it is not possible to define time without reference to frequency. This is something that becomes very apparent, for instance, when we try to arrive at a notion of instantaneous frequency, which the work on frequency modulation highlighted (cf. [1]). Parallel to this, developments in Quantum Physics raised similar questions. In fact, one of the first attempts at tackling the problem of a dual representation appeared in E. Wigner’s 1932 work [15] in that field. A similar approach was independently developed by J. Ville in 1948 within the scope of musical signals, defining [14, p.72]
(59) $F(t,f)=\frac{1}{2\pi}\int\Psi^{*}\left(t-\frac{\nu}{4\pi}\right)\Psi\left(t+\frac{\nu}{4\pi}\right)e^{-j\nu f}dt,$
where the time domain is represented by the complex function $\Psi(t)$24. As can be seen we now have a function of both (radian) frequency $f$ and time $t$.
From another angle, D. Gabor follows a route from De Morgan’s function segmentation. He proposes the use of a Gaussian envelope as a means of focusing the analysis at certain points in time [6, p.435],
(60) $\psi(t)=e^{-\alpha^{2}(t-t_{0})^{2}}\operatorname{cis}(2\pi f_{0}t+\phi),$
and utilising a complex sinusoid representation25 instead of a real signal to facilitate the analysis. This produces a spectral component which is also subject to an envelope, this time in frequency, as its dual,
(61) $\phi(f)=e^{-\left(\frac{\pi}{\alpha}\right)^{2}(f-f_{0})}\operatorname{cis}[-2\pi t_{0}(f-f_{0})+\phi]$
The pair can be used then to represent what Gabor termed acoustical quanta [7]. With this he was able to define exactly the trade-off between time and frequency precision in the analysis. This approach led to the development of the short-time Fourier transform and various very successful time-frequency musical signal processing techniques based on it. Both Wigner-Ville and Gabor opened a new chapter in the story of the Fourier theory, with a lasting impact, but that unfortunately is beyond the scope of this paper.
Notes